Given the 16-bit binary string below for your starting point only, determine its value as requested.

Starting point 1: 0101 0110 1101 1000
Answer #1
Start at the rightmost digit. Take that digit and multiply with 20 (20 = 1). Multiple second digit with 21, third with 22...
0      *  2^0    =  0  *  1  =  0
0     *  2^1    =  0  *  2  =  0
0     *  2^2    =  0  *  4  =  0
1     *  2^3    =  1  *  8  =  8
1     *  2^4    =  1  *  16 =  16
0     *  2^5    =  0  *  32 =  0
1     *  2^6    =  1  *  64 =  64
1     *  2^7    =  1  *  128    =  128
0     *  2^8    =  0  *  256    =  0
1     *  2^9    =  1  *  512    =  512
1     *  2^10   =  1  *  1024   =  1024
0     *  2^11   =  0  *  2048   =  0
1     *  2^12   =  1  *  4096   =  4096
0     *  2^13   =  0  *  8192   =  0
1     *  2^14   =  1  *  16384  =  16384
0     *  2^15   =  0  *  16384  =  0

Add together all products
= 0 + 0 + 0 + 8 + 16 + 0 + 64 + 128 + 0 + 512 + 1024 + 0 + 4096 + 0 + 16384 + 0
= 22232

So, decimal value of 0101 0110 1101 1000 is 22232